pythagoras=25^2
sides= (x+3)
4(x+2)

Answer by stanbon(75887)

pythagoras=25^2
sides= (x+3)
4(x+2)
---------------------
(x+3)^2 + (4(x+2))^2 = 25^2
----
x^2 + 6x + 9 + 16(x^2+4x+4) = 625
---
17x^2 + 70x +73 = 625
----
17x^2 + 70x - 552 = 0
---
x = 4
==========
side: x+3 = 7
side: 4(x+2) = 24
========================